# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

### v1

@@ 3.6. Comparison for Diverse Material Defect Detection Tasks
In material defect detection, the use of different learning methodologies is crucial. Each methodology is designed for specific scenarios and has its own advantages and disadvantages. These are summarized in Table 4, which provides a comparative analysis of the different approaches.
Unsupervised learning is a practical choice when there is a shortage of labelled training data. This approach relies on the algorithm’s ability to identify patterns and anomalies in the data without explicitly labelled examples. The main advantage of unsupervised learning is its flexibility in handling real-world scenarios where obtaining labels for defective and non-defective samples is difficult. This may be due to high costs, time constraints, or the impracticality of manual labelling in complex or nuanced cases. However, this approach is not without limitations. The absence of labelled data, while offering a degree of flexibility, also introduces a potential shortfall in performance metrics. Without explicit labels to guide the learning process, unsupervised algorithms may struggle to achieve the level of precision and accuracy that is attainable with supervised methods. The performance of unsupervised learning algorithms is inherently contingent upon their ability to infer the underlying structure of the data and distinguish between normal variations and genuine defects. This task can be particularly daunting in the context of material defect detection, where the distinction between a defect and a non-defect can be subtle and highly context-dependent.
Supervised learning plays an important role in material defect detection, especially in situations where abundant labelled data is available. It is highly regarded for its effectiveness in achieving high levels of detection accuracy, as well as precise defect localization and classification. The key to its proficiency lies in its utilization of labelled datasets, where each example is annotated with the correct output. This enables the algorithm to learn the complex relationships between input features and desired output labels. One of the strengths of supervised learning is its robustness and predictability. By training on a comprehensive dataset that adequately represents the characteristics of both defects and non-defects, supervised learning models can develop a nuanced understanding of the features that differentiate defects from normal variations in materials. This enables precise defect identification, including size, shape, and location. In turn, this not only facilitates defect detection but also provides valuable insights into their nature and severity. However, the effectiveness of supervised learning is contingent upon the availability and quality of labelled data. Obtaining such data can be labour-intensive and costly, requiring expert knowledge to accurately identify and label defects across potentially large datasets. This is particularly challenging in industrial contexts, where defects can vary greatly and the materials themselves can present a complex range of features that must be correctly interpreted. Furthermore, the dynamic nature of industrial environments and the continuous evolution of materials and manufacturing processes often necessitate frequent updates to the training datasets in order to maintain the accuracy and relevance of supervised learning models. The ongoing need for fresh and accurately labelled data can pose significant logistical and financial challenges, potentially limiting the scalability and adaptability of supervised learning approaches in rapidly changing or resource-constrained settings.
Semi-supervised learning emerges as a particularly effective strategy in situations where labelled data is scarce or expensive to obtain. This approach adeptly bridges the gap between supervised and unsupervised learning by leveraging a small amount of labelled data in conjunction with a larger pool of unlabeled data. The underlying premise is that the unlabeled data, despite the absence of explicit annotations, contains valuable information about the structure and distribution of the data that can be exploited to improve the learning algorithm’s performance. In the context of material defect detection, semi-supervised learning can significantly enhance the model’s ability to identify and classify defects with fewer labelled examples. This is achieved by utilizing the unlabeled data to better understand the feature space’s overall structure, aiding in generalization from the limited labelled samples. The approach can lead to more robust models that are capable of handling the variability and complexity often encountered in material surfaces, without the need for extensive annotated datasets.
Reinforcement learning, on the other hand, presents a paradigm fundamentally different from the traditional supervised or unsupervised learning frameworks. It is particularly suited to interactive and dynamic environments where an agent learns to make decisions by performing actions and receiving feedback in the form of rewards or penalties. This feedback loop enables the agent to learn from its interactions with the environment and optimize its behaviour towards achieving specific goals. In autonomous robotic inspection for defect detection, reinforcement learning can be invaluable. It allows the system to improve its inspection strategies based on the outcomes of its actions, potentially leading to more efficient and effective defect detection processes. However, the training process can be considerably time-consuming and resource-intensive as it requires the agent to explore a vast space of possible actions and their consequences. This exploration process makes it less feasible in scenarios where rapid deployment or real-time decision-making is critical.
Generative learning models, particularly those based on generative adversarial networks (GANs) and variational autoencoders (VAEs), excel in synthesizing realistic defect data, which can significantly enhance the robustness and generalizability of learning algorithms. By generating synthetic data that mimics real-world defects, generative models can augment existing datasets, especially in scenarios where the collection of extensive real defect samples is impractical or impossible due to the rarity or the destructive nature of the defects. However, the deployment of generative learning models is not without its challenges. The process of training these models, particularly GANs, involves a delicate balance between the generator and discriminator components, which requires careful tuning of the model architecture and training parameters. Moreover, generative models are often computationally intensive, requiring significant resources for training and optimization. This can pose a barrier to their application, particularly in resource-constrained environments or scenarios requiring rapid model development and deployment.
@@ 4. Techniques for Composite Material Defect Detection
Composite materials are widely used in high-tech fields such as aerospace, military, etc., Due to its advantages such as lightweight, high strength, and corrosion resistance. However, compared to traditional metal materials, composite materials are more prone to various types of defects, as shown in Table 5, which can lead to a decrease in material performance or even failure. Therefore, it is crucial to conduct defect detection on composite materials.
@@
Machine learning techniques for defect detection in composite materials differ significantly from those used for other types of materials. Composite materials require models capable of handling heterogeneous and complex data due to varied material properties and defect types, requiring advanced feature extraction and more complex models, such as deep learning architectures. Multi-modal data integration is often essential, combining information from various sensors, whereas other materials typically rely on single-mode data. The need for diverse and extensive training datasets is more pronounced in composite materials, which often encounter imbalanced datasets and demand techniques like synthetic data generation. Algorithms for composite materials must be highly robust and adaptable to variations in defect presentation and environmental conditions, and capable of real-time processing. Additionally, the complexity of these models poses challenges for interpretability, increasing the need for techniques to explain model decisions. Transfer learning and domain adaptation are also more crucial for composite materials to apply knowledge from other contexts, whereas simpler models and standard training techniques are often sufficient for other materials.
@@
Non-destructive testing (NDT) technology is one of the key methods for evaluating defects in composite materials. Due to the heterogeneity and anisotropy of composite materials, traditional metal material testing techniques are not always applicable. Common defect detection techniques for composite materials include ultrasonic testing, X-ray testing, thermographic imaging, electronic speckle pattern interferometry, acoustic emission testing, and optical non-destructive testing. The following discussion will delve deeper into these methods.
Ultrasonic Testing is commonly used as one of the significant methods to realize composite defect detection. (mcknight20243, ) develops a deep learning approach with 3-D convolutional neural networks (CNNs) for detecting defects in carbon fibre-reinforced polymer composites using volumetric ultrasonic data. To manage the high costs of experimental data, the study uses synthetically generated volumetric data, enhancing the model’s ability to utilize spatial and temporal information. Three architectures were tested: a standard model, a custom model with cuboidal kernels suited for ultrasonic data, and an optimized model developed through neural architecture search (NAS). The NAS-derived model performed best, achieving a significant 22.4% improvement in accuracy over other models and processing data in under 0.5 seconds, proving highly efficient for real-time defect detection. Then, (lu2024research, ) discusses enhancing ultrasonic testing for defect detection in industrial applications, highlighting its current focus on surface scanning and internal defect classification. The study utilizes finite element simulation software to detect simple defects in specimens, with accuracy validated through ultrasonic experiments. Features are extracted using signal processing methods, and a neural network with a Gaussian weight fusion model is used to visualize results. The approach extends to training on signals from simple defects to detect and image complex defects, aiming to improve the accuracy of imaging internal defect morphology. Meanwhile, (oztacs2024image, ) examines the detection of internal damages in composite materials, crucial for industries like aircraft manufacturing. Using glass and carbon fibre-reinforced epoxy composites with artificially induced damages, the materials were subjected to impact tests and analyzed with ultrasonic scanning. The results revealed significant differences in damaged areas, especially in glass fibre-reinforced composites, where the damaged area was up to five times larger than in undamaged samples. Carbon fibre composites with artificial damages showed perforations from impact tests, unlike undamaged samples. Image processing algorithms effectively detected and quantified these barely visible damages, demonstrating the effectiveness of the applied non-destructive testing methods. Besides, X-ray Testing also plays an important role in performing the composite material defect detection. (gong2020deep, ) propose the use of deep transfer learning to enhance the detection of defects in X-ray images of aeronautic composite materials. This approach addresses the limitations of traditional non-destructive testing methods, which are often manual and subjective. By leveraging a model trained on limited samples, the authors achieve 96% classification accuracy, demonstrating the model’s effectiveness in automatic defect detection.
Table 5. Common Composite Material Defects.
Defect Type Defect Manifestation Defect Formation Cause
delamination (de2024understanding, ; liu2020data, )
separation of layers
within the composite material
inadequate bonding during manufacturing
or by mechanical impact
porosity (habibi2021effect, ; al2020effects, )
presence of voids or air pockets
within the material
improper curing, resin infusion, or insufficient
vacuum pressure during manufacturing.
fiber misalignment (zhang2021fibre, )
misalignment or disorientation of reinforcing
fibers within the composite
improper layup or handling
during manufacturing
fiber breakage (suriani2021delamination, ; zhang2021fibre, )
breakage or damage to
individual reinforcing fibers
excessive mechanical stress or impact
matrix cracking (fu2022review, ; mardanshahi2020detection, )
cracks or fractures within the resin
matrix of the composite material
thermal cycling, mechanical loading, or
environmental factors such as moisture ingress
Impact damage (zimmermann2020review, ) surface or subsurface damage impact from foreign objects
inclusions (gong2020deep, ; roy2021compound, )
foreign particles or contaminants
embedded within the composite material
contaminants or impurities introduced during
the melting and solidification processes
fiber pull-out (hemmatian2023machine, ; de2021novel, )
Partial or complete removal of reinforcing
fibers from the matrix
inadequate bonding or insufficient
fiber-matrix adhesion
Another detection technical approach is based on thermographic Imaging for composite material defects. (liu2024multi, ) presents a new vision transformer model for detecting and quantifying defects in composites using thermography. It improves data handling with a novel sampling strategy for infrared data and uses multiple-feature fusion to enhance defect detection. The model, leveraging self-attention mechanisms, excels in accuracy and generalization, effectively identifying defects across composite images. (zhang2024one, ) introduces a novel non-destructive testing method, one-dimensional deep convolutional autoencoder active infrared thermography (1D-DCAE-AIRT), for enhancing defect detection in fibre-reinforced polymer composites. Traditional active infrared thermography methods face challenges with uneven heating, affecting defect visibility. The 1D-DCAE-AIRT method preprocesses thermal images, extracts high-level features using a deep convolutional autoencoder, and uses these features to improve the visualization of defects. Tests on composite specimens with artificially introduced defects showed that this method surpasses traditional techniques like FFT, PCA, ICA, and PLSR in revealing internal defects. (deng2024attention, ) introduces a spatiotemporal-based deep learning model with an attention mechanism designed to analyze thermograms more effectively. The model processes thermal image sequences to characterize barely visible impact damages (BVID) in composite materials, linking these damages to specific impact energy levels. Validated with 100 composite specimens subjected to varying impact forces, the model achieves a classification accuracy of over 95%. This approach offers significant potential for Structural Health Monitoring (SHM), particularly in the aviation industry, by facilitating targeted predictive maintenance and enhancing safety through accurate impact damage analysis.
Optical non-destructive testing has been a promising way to conduct effective and efficient defect detection for composite materials. (palka2024automatic, ) introduces a method using terahertz time-domain spectroscopy (TDS) to automatically detect defects in glass fibre-reinforced polymer (GFRP). Employing a high-speed TDS scanner with an advanced sampling system, the method successfully detected 87% of inclusions and 100% of delaminations in GFRP specimens. A thresholding algorithm based on Gaussian fitting of histograms helped pre-select potential defects, also providing depth information. The results highlight the system’s effectiveness in defect detection in GFRP. (mahmoud2024enhancing, ) proposes a method using hyperspectral imaging (HSI) to detect defects in carbon fibre composites (CFC), which are crucial in various industries for their strength and stiffness. By employing HSI, the method captures detailed spectral data that highlights differences in reflectance between defective and non-defective areas, particularly in the 440 nm to 600 nm range. Integrating HSI with k-means clustering automates defect characterization, enhancing detection speed and accuracy. The approach uses specific wavelengths as markers to measure fracture depths and produces 3D images of defects, offering a precise and efficient tool for evaluating CFC components in industrial applications. (kazemi2024preliminary, ) presents a prototype intelligent quality inspection tool that uses camera vision to monitor fibre distortion in fabric composites during forming processes. Utilizing data from about 30 samples and over 1200 images, the tool employs machine learning algorithms, specifically K-nearest neighbours and Support Vector Machine, to identify defects. These models achieved an error rate of less than 5% in differentiating between defective and non-defective parts, proving effective even amid background noise and image blurriness. This AI tool can be used in a smart factory by operators or robots to scan and assess formed parts efficiently.
More detection techniques include acoustic emission testing and electronic speckle pattern interferometry. (lan2024impact, ) details a method using local defect resonance (LDR) to detect barely visible impact damages (BVIDs) in woven carbon fibre-reinforced polymer (CFRP) laminates through laser-ultrasonic guided waves. The technique involves exciting broadband guided waves with an Nd: YAG pulsed laser, directed using a galvano mirror system, while a fixed broadband piezoelectric sensor records the waves. The recorded data are then processed to visualize the wavefield and extract LDR frequencies, producing high-contrast images of BVIDs. Validated on a CFRP laminate with a single BVID, the method outperformed traditional root mean square (RMS) techniques and successfully demonstrated the detection of multiple BVIDs. (tu2024infrared, ) presents an infrared-induced laser shearography (IILS) method that improves defect detection in composite structures by capturing dynamic deformation data. Utilizing the temperature-deformation connection theory, the method investigates the relationship between dynamic deformations and defects through simulations and experiments. Feature extraction algorithms, particularly tensor robust principal component analysis (TRPCA), are employed to denoise and enhance imaging. This method significantly improves image quality and detection capabilities, validated by metrics such as two-dimensional entropy, contrast, and signal-to-noise ratio. Overall, IILS enhances the reliability and stability of non-destructive testing in industrial applications. (yoon2023defect, ) introduce a non-destructive technique using deep learning to detect delamination in laminated AS4/PEEK composites. The approach utilizes highly nonlinear solitary waves (HNSWs) captured by a granular crystal sensor, which feeds data into an optimized convolutional neural network (CNN). By adjusting CNN parameters and incorporating multiple pre-processed signal modes, the study achieves over 90% accuracy in detecting defects, showcasing the potential for real-time, accurate defect detection in composites.

### v3

@@ 3.6. Comparison for Diverse Material Defect Detection Tasks
Performance Comparison Among Different ML Techniques
In MDD, the use of different learning methodologies is crucial. Each methodology is designed for specific scenarios and has its own advantages and disadvantages. These are summarized in Table 6, which provides a comparative analysis of the different approaches. The choice of learning technique in MDD depends heavily on data availability, computational resources, and the complexity of the application context. UL provides flexibility in scenarios where labeled data is scarce, but its lack of explicit guidance often limits precision (kumar2021histogram, ; yuan2015improved, ). SL, leveraging methods like RCNN and YOLO, achieves high accuracy and precise defect localization but is constrained by the need for extensive labeled datasets and frequent updates to maintain relevance (tabernik2020segmentation, ). SSL bridges this gap by utilizing small labeled datasets alongside larger unlabeled ones, offering a cost-effective solution for complex and variable defect scenarios (chu2020neural, ). RL excels in dynamic environments through feedback-driven optimization, enabling advanced defect detection strategies but requiring significant computational resources and long training times (chu2020neural, ; chung2022reinforcement, ). Meanwhile, GL models, such as GANs and VAEs, enhance dataset diversity by synthesizing realistic defect samples, improving robustness and generalizability in data-scarce environments, though they are computationally intensive and require careful tuning (zhang2021defect, ; jain2022synthetic, ). Ultimately, while SL and SSL dominate in achieving accuracy and generalizability, RL and GL present innovative solutions for dynamic and data-augmented scenarios, balancing strengths and limitations to meet diverse industrial needs.
Table 7. Performance comparison of traditional machine learning and deep learning techniques for material defect detection.
Criteria Traditional ML Methods Deep Learning Methods
Accuracy Moderate accuracy, heavily dependent on handcrafted features. High accuracy due to automatic feature extraction and end-to-end learning.
Efficiency Faster training and inference with low computational requirements. Computationally intensive, requiring significant hardware resources for training.
Robustness Limited robustness to variations in defect types or environmental noise. High robustness to defect variations, but may require large, diverse datasets for generalization.
Feature Handling Relies on manual feature engineering, which can be time-consuming and domain-specific. Automatically extracts hierarchical features, reducing dependency on manual effort.
Scalability Effective for small datasets but struggles with scalability for complex problems. Scales well to large datasets and complex defect detection tasks.
Applicability Suitable for simpler, well-defined problems with structured data. Excels in complex, high-dimensional problems with diverse data modalities.
Data Requirements Performs well with limited labeled data. Requires extensive labeled data for training to achieve optimal results.
Comparative Analysis of Traditional ML and DL Approaches
Traditional ML methods, such as feature statistical techniques and filter-based approaches, have been widely applied in material defect detection due to their efficiency and simplicity. Techniques like the GLCM and Gabor filters excel in detecting well-defined defect types, offering fast processing with minimal computational demands. However, these methods rely on manually engineered features, which limit their adaptability to complex or heterogeneous defects, reducing their robustness in real-world scenarios.
In contrast, DL methods, including CNNs and Vision Transformers (ViTs), demonstrate significant improvements in accuracy and adaptability, particularly for intricate and overlapping defect types. By automatically extracting hierarchical features, DL models excel in diverse and challenging defect patterns. For example, CNNs are effective in detecting subtle surface defects, while ViTs capture global contextual information in high-resolution defect images (liu2024multi, ). Nevertheless, DL methods require substantial computational resources and large annotated datasets, making traditional ML approaches more practical in resource-constrained or simpler detection environments. Table 7 summarizes these differences, highlighting the trade-offs between accuracy, efficiency, and robustness for both approaches.
Performance Comparison On Real-world Applications
MDD is a cornerstone of industrial quality assurance, where ML techniques have proven to be transformative. Real-world applications, such as steel surface defect detection, fabric inspection, and concrete crack analysis, not only test the robustness of these techniques but also provide a foundation for benchmarking advancements. Across these applications, researchers have leveraged diverse datasets, such as NEU (song2014automatic, ), GC10-Det (Lv2020, ), MVTec AD (bergmann2019mvtec, ), and NanoTWICE (Carrera2017, ), to ensure standardized and fair evaluation. These benchmarks enable a direct comparison of traditional and modern approaches, highlighting how ML techniques have advanced the state of defect detection.
As shown in Table 8, advanced methods such as ViTs and YOLOv8 consistently outperform traditional techniques, achieving metrics like mAP scores above 98% in steel surface (song2014automatic, ) and textured surface defect detection (bergmann2019mvtec, ). Conversely, classical methods, such as histogram-based thresholding, exhibit limitations in handling complex scenarios, with lower F1-scores observed on datasets like SDNET 2018 (Dorafshan2018, ). This performance gap underscores the critical role of DL in adapting to diverse defect types. The adoption of such methods has not only improved accuracy and reliability but also established a roadmap for future innovations in industrial defect detection, ensuring methods are scalable and applicable across a wide range of domains.
Table 8. Quantitative performance comparison on MMD real-world applications
Real-world Application Ref. Method Dataset (Ref.) Evaluation Metric and Score
Steel Surface Defect Detection (song2014automatic, ) Scattering Convolution Network NEU (song2014automatic, ) mAP: 96.3%
(cohn2021unsupervised, ) K-Means Clustering GC10-Det (Lv2020, ) Clustering Score: 91%
(wang2024, ) Improved YOLOv8 NEU-DET (wang2024, ) mAP: 98.1%
(zhu2024mrp, ) MRP-YOLO (Modified YOLOv8) NEU-DET (wang2024, ) mAP: 95.4%
Fabric Defect Detection (SilvestreBlanes2019, ) GLCM + SVM AITEX (SilvestreBlanes2019, ) F1-Score: 87%
(zhu2015autocorrelation, ) Autocorrelation + GLCM Tianchi Cloth (Qi2023, ) Precision: 88.5%
Concrete Crack Detection (shalaby2024condition, ) Histogram-based Threshold Detection SDNET 2018 (Dorafshan2018, ) F1-Score: 74.11%
(Li2019, ) Deep CNN for Crack Detection Surface Crack (Ozgenel2018, ) mAP: 94.6%
(Dorafshan2018, ) Deep CNN SDNET 2018 (Dorafshan2018, ) mAP: 92.3%
Texture Surface Analysis (bergmann2019mvtec, ) Unsupervised Anomaly Detection MVTec AD (bergmann2019mvtec, ) AUC: 95.4%
(li2021anomaly, ) Self-organizing Map for Anomaly Detection (SOMAD) MVTec AD (bergmann2019mvtec, ) mAP: 97.9%
(tao2023vitalnet, ) Vision Transformer MVTec AD (bergmann2019mvtec, ) mAP: 99.0%
(Lv2020, ) Deep Learning-Based Feature Extraction Kylberg (Kylberg2011, ) Recall: 92.7%
Industrial Surface Defects (Tabernik2020, ) Segmentation-Based DL Method Kolektor (Tabernik2020, ) Mean IoU: 93%
(Tang2019, ) Group Pyramid Pooling (GPP) DeepPCB (Tang2019, ) Precision: 90.8%
Miscellaneous Applications (olimov2022uzadl, ) Unsupervised Laplacian-based SOM NanoTWICE (Carrera2017, ) mAP: 89.3%
(tao2023vitalnet, ) Vision Transformer NanoTWICE (Carrera2017, ) AUC: 94.5%
Emerging Optimization Techniques for Real-Time Detection
Recent advancements in neural architecture search (NAS) and model compression techniques have significantly improved the performance of ML models for real-time defect detection. NAS enables automated network design optimized for accuracy, latency, and resource constraints, making it suitable for industrial applications with stringent time requirements (elsken2019neural, ; liu2021survey, ). Additionally, techniques such as knowledge distillation have been effectively used to transfer the performance of large models to lightweight architectures. These smaller models, as demonstrated by Hinton et al. (hinton2015distilling, ), retain high accuracy while ensuring low latency for time-critical tasks.
Furthermore, hybrid edge-cloud computing frameworks have emerged as practical solutions for achieving high throughput and low latency in industrial settings. By processing time-sensitive tasks on edge devices and offloading computationally intensive operations to the cloud, these frameworks balance speed and scalability, as shown in steel surface and fabric defect detection systems (xu2020edge, ; kang2017neurosurgeon, ). Combined with hardware-specific optimizations and real-time processing pipelines, these advancements demonstrate the growing feasibility of deploying DL models in industrial environments where latency and throughput are critical.
@@ 4. Techniques for Composite Material Defect Detection
Composite materials (CMs), widely used in high-tech fields such as aerospace and military due to their advantages of lightweight, high strength, and corrosion resistance, present unique challenges for defect detection. Unlike traditional homogeneous materials like metals, composites are anisotropic and heterogeneous, making them more prone to various types of defects, which can significantly degrade material performance or even lead to failure. Addressing these challenges requires advanced nondestructive evaluation (NDE) techniques and tailored ML models. While many general MDD algorithms discussed earlier are applicable, this section emphasizes adaptations and innovations specifically designed for CMs, including customized feature extraction methods and learning frameworks.
4.1. Overview of Challenges in Composite Materials
@@
CMs require models capable of handling heterogeneous and complex data due to varied material properties and defect types, requiring advanced feature extraction and more complex models, such as DL architectures. Multi-modal data integration is often essential, combining information from various sensors, whereas other materials typically rely on single-mode data. Additionally, algorithms for CMs must be highly robust and adaptable to variations in defect presentation and environmental conditions. The inherent texture complexity in composites also emphasizes the need for texture analysis methods such as Gabor filters and GLCM, discussed later in this section.
4.2. Texture Feature Analysis in Composite Materials
Texture features play a vital role in detecting defects in composites due to their heterogeneous nature. Established methods like Gabor filters and GLCM have been widely utilized to extract and analyze texture features in composite defect detection (osman2020automatic, ; hui2019application, ; li2018gabor, ; sharma2021grey, ; vasan2019defect, ). Gabor filters are effective in capturing spatial frequency, orientation, and texture information, making them suitable for identifying localized surface defects in composites (li2018gabor, ). By applying multi-scale and multi-orientation Gabor filters, researchers have successfully extracted texture features that highlight defect patterns against complex composite backgrounds. GLCM, on the other hand, provides a statistical approach to analyzing texture by evaluating the spatial relationship of pixel intensities. Features derived from GLCM, such as contrast, correlation, energy, and homogeneity, offer a robust representation of composite material textures, enabling effective detection of defects like delamination or matrix cracking (sharma2021grey, ; vasan2019defect, ). Together, these methods enhance the precision of detecting defects in composites, addressing challenges posed by the materials’ anisotropic and heterogeneous properties.
4.3. Non-Destructive Testing Techniques
@@
NDT technology is one of the key methods for evaluating defects in CMs. Due to the heterogeneity and anisotropy of CMs, traditional metal material testing techniques are not always applicable. Common defect detection techniques for CMs include ultrasonic testing, X-ray testing, thermographic imaging, electronic speckle pattern interferometry, acoustic emission testing, and optical non-destructive testing. Below, we delve deeper into these methods based on their application domains.
4.3.1. Ultrasonic Testing
Ultrasonic testing is a widely used method for composite defect detection due to its ability to penetrate deep into materials and detect internal defects. For instance, (mcknight20243, ) develops a DL approach using 3-D CNNs for detecting defects in carbon fibre-reinforced polymer composites using volumetric ultrasonic data. Meanwhile, (oztacs2024image, ) examines the detection of internal damages in CMs using ultrasonic scanning, emphasizing its effectiveness in identifying defects like delamination. (lu2024research, ) discusses enhancing ultrasonic testing with advanced signal processing techniques, which improve the detection of complex internal defects.
4.3.2. X-Ray Testing
X-ray testing provides high-resolution imaging for detecting internal defects in CMs. (gong2020deep, ) proposes the use of deep transfer learning to enhance the detection of defects in X-ray images of aeronautic CMs, achieving a classification accuracy of 96%.
4.3.3. Thermographic Imaging
Thermographic imaging has emerged as a powerful tool for detecting surface and subsurface defects in CMs. (liu2024multi, ) presents a vision transformer model for thermography-based defect detection, while (zhang2024one, ) introduces a one-dimensional deep CAE active infrared thermography (1D-DCAE-AIRT) approach to enhance defect visualization. Furthermore, (deng2024attention, ) leverages a spatiotemporal-based DL model with an attention mechanism to analyze thermograms for detecting barely visible impact damages (BVID).
4.3.4. Optical Non-Destructive Testing
Optical methods are another promising approach for composite MDD. (palka2024automatic, ) introduces a method using terahertz time-domain spectroscopy (TDS) to automatically detect defects in glass fibre-reinforced polymer (GFRP). Meanwhile, (mahmoud2024enhancing, ) employs hyperspectral imaging (HSI) to detect defects in carbon fibre composites (CFC), enhancing detection speed and accuracy through spectral data analysis.
4.3.5. Visualization for Composite Defects
Visualization plays a critical role in composite MDD by providing interpretable representations of defects captured through various imaging techniques. As presented in (Vavilov2020, ), multiple vision systems are employed to generate diverse defect visualizations. Optically-stimulated infrared thermography with different heating sources (tube or lamp heating) emphasizes surface and subsurface defects by creating high-contrast thermal maps. Ultrasonic techniques, including laser vibrometry and laser ultrasonics, enable precise imaging of internal defect structures by utilizing wave-based properties. These methods offer complementary perspectives on defect characterization, ranging from pixel-level defect identification to volumetric internal defect mapping. The integration of such visualizations into ML frameworks significantly enhances the interpretability and reliability of defect detection systems, providing a robust foundation for evaluating the structural integrity of CMs.
4.4. Multimodal Data Integration for Composite Material Defect Detection
The integration of multimodal data has emerged as a powerful approach for enhancing the accuracy and robustness of defect detection in CMs. By combining complementary data from different sources such as thermal imaging, ultrasonic testing, and optical sensing, multimodal integration addresses the limitations of single-modal methods, such as insufficient feature representation and sensitivity to specific environmental conditions. Recent advancements in this area have demonstrated the effectiveness of integrating data from multiple modalities to provide a more comprehensive understanding of defects in composite structures.
For instance, the fusion of phased-array ultrasonics (PAUT) and pulsed thermography (PT) has been successfully implemented using an Automatic Defect Recognition (M-ADR) system, enabling detailed structural evaluations of CMs. This multimodal approach not only improves the detection of internal defects like delamination but also enhances the resolution and reliability of surface defect assessments (sudharsan2024multi, ). Similarly, machine vision transformer networks (M-VIT-DDQ) have been developed to process multimodal thermal and optical data for defect detection in composites, leveraging multi-feature fusion to detect delamination and matrix cracking with high accuracy (liu2024multi, ). These techniques demonstrate the potential of combining multiple sensing modalities to achieve more accurate and versatile defect detection.
In addition to thermal and ultrasonic modalities, the synchronous fusion of optical and electromagnetic imaging has been explored to detect overlapping defects in complex composite geometries. By aligning and combining data from these modalities, researchers have achieved significant improvements in detecting subsurface defects that are challenging to identify using individual techniques (ye2023synchronous, ). Advanced ML models, such as DL frameworks and transformers, play a critical role in processing and interpreting multimodal inputs, enabling efficient feature extraction and integration of heterogeneous data.
Despite these advancements, challenges remain in the synchronization, dimensionality reduction, and noise suppression of multimodal data. To address these issues, recent studies have explored innovative solutions, such as tensor-based data analytics and DL-based fusion frameworks, which optimize the processing and integration of multimodal information. Moving forward, the integration of real-time processing capabilities and the development of domain-adaptive multimodal frameworks will be critical for advancing defect detection in CMs. By leveraging the strengths of multimodal data integration, researchers and practitioners can overcome the limitations of traditional single-modal approaches, achieving improved accuracy, efficiency, and reliability in composite MDD.